# Stochastic Gradient Langevin Unlearning

### v1

@@ Appendix C Existence of limiting distribution
    Strong aperiodicity. Since $\mathcal{C}_{R}$ and $\bigcup_{i\in\mathcal{B}^{n/b}}\nabla f(\mathcal{C}_{R};\mathbf{d}_{i})$ are bounded and the density of $W_{t}$ has a uniform positive lower bound on any bounded domain, there exists a non-zero multiple of the Lebesgue measure, say $\nu_{1}$, satisfying that
    $\mathbb{P}(x_{1}\in A\ |\ x_{0}=x)\geq\nu_{1}(A),\quad\forall~{}x\in\mathcal{C}_{R},\ A\in\mathcal{B}(\mathcal{C}_{R}).$
    Then $\{x_{t}\}$ is strongly aperiodic by the equation above and $\nu_{1}(\mathcal{C}_{R})>0$ (see Section 5.4.3 in (Meyn & Tweedie, 2012) for definition).
@@
Theorem C.2.
Under the same assumptions as in Proposition C.1, the Markov chain $\{x_{t}\}$ admits a unique invariant probability measure $\nu_{\mathcal{D}|\mathcal{B}}$ on $\mathcal{B}(\mathcal{C}_{R})$. Furthermore, for any $x\in\mathcal{C}_{R}$, the distribution of $x_{t}=x_{t}^{0}$ generated by the learning process in Algorithm 1 conditioned on $x_{0}=x$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ as $t\to\infty$.
@@
It has been proved in Proposition C.1 that $\{x_{t}\}$ is strongly aperiodic and recurrent with an invariant measure. Consider any $A\in\mathcal{B}(\mathcal{C}_{R})$ with $\psi(A)>0$ and use the same settings and notations as in the proof of Proposition C.1. There exists $T\geq 0$, $\delta>0$, and $B\in\mathcal{B}(\mathcal{C}_{R})$ with $\textrm{Leb}(B)>0$, such that $\mathbb{P}(x_{T}\in A\ |\ x_{0}=x)\geq\delta,\ \forall~{}x\in B$. This implies that for any $t\geq 0$ and any $x\in\mathcal{C}_{R}$,
$\mathbb{P}(x_{t+T+1}\in A\ |\ x_{t}=x)=\mathbb{P}(x_{T+1}\in A\ |\ x_{0}=x)\geq\mathbb{P}(x_{T+1}\in A\ |\ x_{1}\in B,x_{0}=x)\cdot\mathbb{P}(x_{1}\in B\ |\ x_{0}=x)\geq\epsilon,$
@@
$\epsilon=\delta\cdot\inf_{y\in\mathcal{C}_{R}}\mathbb{P}(x_{1}\in B\ |\ x_{0}=y)>0,$
@@
This verifies that the chain $\{x_{t}\}$ is Harris recurrent (see Section 9 in (Meyn & Tweedie, 2012) for definition). It can be further derived that for any $x\in\mathcal{C}_{R}$,
@@
The bound above is uniform for all $x\in\mathcal{C}_{R}$ and this implies that $\mathcal{C}_{R}$ is a regular set of $\{x_{t}\}$ (see Section 11 in (Meyn & Tweedie, 2012) for definition). Finally, one can apply Theorem 13.0.1 in (Meyn & Tweedie, 2012) to conclude that there exists a unique invariant probability measure $\nu_{\mathcal{D}}$ on $\mathcal{B}(\mathcal{C}_{R})$ and that the distribution of $x_{t}$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ conditioned on $x_{0}=x$ for any $x\in\mathcal{C}_{R}$. ∎
Appendix D Proof of Theorem 3.2
@@
Definition D.1 (Shifted Rényi divergence).
@@
Lemma D.2 (Data-processing inequality for Rényi divergence, Lemma 2.6 in (Altschuler & Talwar, 2022b)).
@@
$D_{\alpha}(h_{\#}\mu||h_{\#}\nu)\leq D_{\alpha}(\mu||\nu).$ (6)
Proposition D.3 (Weak Triangle Inequality of Rényi divergence, Corollary 4 in (Mironov, 2017)).
@@
Recall that from the sketch of proof, we have defined the six distributions: $\nu_{\mathcal{D}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0},\rho_{K|\mathcal{B}}^{0}$ are the stationary distribution of the learning process, distribution at epoch $T$ of the learning process and distribution at epoch $K$ of the unlearning process. Note that we learn on dataset $\mathcal{D}$ and fine-tune on $\mathcal{D}^{\prime}$. On the other hand, the corresponding distributions of “adjacent” processes that learn on $\mathcal{D}^{\prime}$ and still unlearn on $\mathcal{D}^{\prime}$ are denoted as $\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime},\rho_{K|\mathcal{B}}^{0,\prime}$ similarly. Note that $\nu_{T|\mathcal{B}}^{0,\prime}$ is the distribution of retraining from scratch on $\mathcal{D}^{\prime}$, and we aim to bound $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})$ for all possible $\mathcal{D},\mathcal{D}^{\prime}$ pairs. By Proposition D.3, we know that for any $\alpha>1$, by choosing $p=q=2$, we have
@@
Recall the Definition 2.1 that $d_{\alpha}(P,Q)=\max(D_{\alpha}(P||Q),D_{\alpha}(Q||P))$ for distributions $P,Q$. The above inequality is correct since consider any distributions $P,Q,R$, by Proposition D.3 we have
@@
Finally, due to data-processing inequality for Rényi divergence (Lemma D.2), applying the final projection step does not increase the corresponding Rényi divergence. As a result, by Lemma 3.4 we have
@@
Appendix E Proof of Theorem 3.11
@@
The proof is a direct application of triangle inequality, Lemma 3.6 and 3.5. We will prove it by induction. For the base case $s=1$ it trivial, since $\bar{\nu}_{\mathcal{D}_{0}}=\nu_{\mathcal{D}_{0}}$ as we choose $y_{0}^{0,(0)}=x_{\infty}^{0}$. Thus by our definition that $Z_{\mathcal{B}}$ is the upper bound of $W_{\infty}(\nu_{\mathcal{D|\mathcal{B}}},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ for any adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$. Apparently, we also have
@@
Appendix F Proof of Corollary 3.8
@@
Appendix G Proof of Corollary 3.9
We first restate the lemma in (Ye & Shokri, 2022), which is an application of the joint convexity of KL divergence.
Lemma G.1 (Lemma 4.1 in Ye & Shokri (2022)).
@@
The proof of Corollary 3.9 directly follows by repeating the proof of Lemma 3.8 but with the $\mathcal{B}$ dependent $Z_{\mathcal{B}}$ described in Lemma 3.5. Then leverage Lemma G.1 leads to the result.
Appendix H Proof of Corollary 3.12
Corollary (Upper bound for $W_{\infty}$ condition on $\mathcal{B}$ for batch unlearning).
Consider the learning process in Algorithm 1 on adjacent dataset $\mathcal{D}$ and $\mathcal{D}^{\prime}$ that differ in $1\leq S$ points and a fixed mini-batch sequence $\mathcal{B}$. Assume $\forall\mathbf{d}\in\mathcal{X}$, $f(x;\mathbf{d})$ is $L$-smooth, $M$-Lipschitz and $m$-strongly convex in $x$. Let the index of different data point between $\mathcal{D},\mathcal{D}^{\prime}$ belongs to mini-batches $\mathcal{B}^{j_{0}},\mathcal{B}^{j_{1}},\ldots,\mathcal{B}^{j_{G-1}}$, each of which contains $1\leq S_{j_{g}}\leq b$ such that $\sum_{g=0}^{G-1}S_{j_{g}}=S$. Then for $\eta\leq\frac{1}{L}$, we have
$W_{\infty}(\nu_{T|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})\leq\min\left(\frac{1-c^{T}}{1-c}\sum_{g=0}^{G-1}c^{n/b-j_{g}-1}\frac{2\eta MS_{j_{g}}}{b},2R\right),$
where $c=(1-\eta m)$.
@@
The proof is a direct generalization of Theorem 3.5. Recall that for the per-iteration bound within an epoch, there are two possible cases: 1) we encounter the mini-batch $\mathcal{B}^{j}$ that contains indices of replaced points 2) or not. This is equivalently to 1) $j\in\{j_{g}\}_{g=0}^{G-1}$ or 2) $j\notin\{j_{g}\}_{g=0}^{G-1}$.
@@
Using this bound and taking infimum over all possible coupling over $\nu_{t|\mathcal{B}}^{j},\nu_{t|\mathcal{B}}^{j,\prime}$, we have
$W_{\infty}(\nu_{t|\mathcal{B}}^{j+1},\nu_{t|\mathcal{B}}^{j+1,\prime})\leq(1-\eta m)W_{\infty}(\nu_{t|\mathcal{B}}^{j},\nu_{t|\mathcal{B}}^{j,\prime})+\frac{2\eta MS_{j_{g}}}{b}\mathbf{1}[j=j_{g}],$ (42)
@@
$W_{\infty}(\nu_{t+1|\mathcal{B}}^{0},\nu_{t+1|\mathcal{B}}^{0,\prime})\leq(1-\eta m)^{n/b}W_{\infty}(\nu_{t|\mathcal{B}}^{0},\nu_{t|\mathcal{B}}^{0,\prime})+\sum_{g=0}^{G-1}(1-\eta m)^{n/b-j_{g}-1}\frac{2\eta MS_{j_{g}}}{b}.$ (43)
The rest analysis is the same as those in Theorem 3.5, where we iterate over $T$ epochs and then simplify the expression using geometric series. Thus we complete the proof. ∎
Appendix I Proof of Lemma 3.5
@@
We proof the bound of one iteration and iterate over that result. Given an mini-batch sequence $\mathcal{B}$, assume the adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$ differ at index $i_{0}\in[n]$ and $i_{0}$ belongs to $j_{0}^{th}$ mini-batch (i.e., $i_{0}\in\mathcal{B}^{j_{0}}$. For simplicity, we drop the condition on $\mathcal{B}$ for all quantities in the proof as long as it is clear that we always condition on a fixed $\mathcal{B}$. Let us denote $\psi_{j}(x)=x-\frac{\eta}{b}\sum_{i\in\mathbf{B}^{j}}\nabla f(x;\mathbf{d}_{i}^{\prime})$ and similar for $\psi_{j}^{\prime}$ on $\mathcal{D}^{\prime}$. Note that for some coupling of $\nu_{t}^{j},\nu_{t}^{j,\prime}$ denoted as $\gamma_{t}^{j}$,
@@
Now, note that by definition the noise $W_{t}^{j}$ is independent of $x_{t}^{j}$, we can simply choose the coupling so that $W_{t}^{j}=W_{t}^{j,\prime}$. So the last term is 00. For the first term, there are two possible cases: 1) $j=j_{0}$ and 2) $j\neq j_{0}$. For case 1), we have
@@
Using this bound and taking infimum over all possible coupling over $\nu_{t|\mathcal{B}}^{j},\nu_{t|\mathcal{B}}^{j,\prime}$, we have
$W_{\infty}(\nu_{t|\mathcal{B}}^{j+1},\nu_{t|\mathcal{B}}^{j+1,\prime})\leq(1-\eta m)W_{\infty}(\nu_{t|\mathcal{B}}^{j},\nu_{t|\mathcal{B}}^{j,\prime})+\frac{2\eta M}{b}\mathbf{1}[j=j_{0}],$ (57)
@@
$W_{\infty}(\nu_{t+1|\mathcal{B}}^{0},\nu_{t+1|\mathcal{B}}^{0,\prime})\leq(1-\eta m)^{n/b}W_{\infty}(\nu_{t|\mathcal{B}}^{0},\nu_{t|\mathcal{B}}^{0,\prime})+(1-\eta m)^{n/b-j_{0}-1}\frac{2\eta M}{b},$ (58)
@@
$W_{\infty}(\nu_{T|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})\leq(1-\eta m)^{Tn/b}W_{\infty}(\nu_{0|\mathcal{B}}^{0},\nu_{0|\mathcal{B}}^{0,\prime})+\sum_{t=0}^{T-1}(1-\eta m)^{(t+1)n/b-j_{0}-1}\frac{2\eta M}{b}$ (59)
@@
since $\nu_{0|\mathcal{B}}^{0}=\nu_{0|\mathcal{B}}^{0,\prime}$ due to the same initialization. Let $c=(1-\eta m)$. Then the bound can be simplified as
@@
Appendix J Proof of Lemma 3.6

### v2

@@ Appendix E Existence of limiting distribution
    Strong aperiodicity. Since $\mathcal{C}$ and $\bigcup_{i\in\mathcal{B}^{n/b}}\nabla f(\mathcal{C};\mathbf{d}_{i})$ are bounded and the density of $W_{t}$ has a uniform positive lower bound on any bounded domain, there exists a non-zero multiple of the Lebesgue measure, say $\nu_{1}$, satisfying that
    $\mathbb{P}(x_{1}\in A\ |\ x_{0}=x)\geq\nu_{1}(A),\quad\forall~{}x\in\mathcal{C},\ A\in\mathcal{B}(\mathcal{C}).$
    Then $\{x_{t}\}$ is strongly aperiodic by the equation above and $\nu_{1}(\mathcal{C})>0$ (see Section 5.4.3 in [23] for definition).
@@
Theorem E.2.
Under the same assumptions as in Proposition E.1, the Markov chain $\{x_{t}\}$ admits a unique invariant probability measure $\nu_{\mathcal{D}|\mathcal{B}}$ on $\mathcal{B}(\mathcal{C})$. Furthermore, for any $x\in\mathcal{C}$, the distribution of $x_{t}=x_{t}^{0}$ generated by the learning process in Algorithm 1 conditioned on $x_{0}=x$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ as $t\to\infty$.
@@
It has been proved in Proposition E.1 that $\{x_{t}\}$ is strongly aperiodic and recurrent with an invariant measure. Consider any $A\in\mathcal{B}(\mathcal{C})$ with $\psi(A)>0$ and use the same settings and notations as in the proof of Proposition E.1. There exists $T\geq 0$, $\delta>0$, and $B\in\mathcal{B}(\mathcal{C})$ with $\textrm{Leb}(B)>0$, such that $\mathbb{P}(x_{T}\in A\ |\ x_{0}=x)\geq\delta,\ \forall~{}x\in B$. This implies that for any $t\geq 0$ and any $x\in\mathcal{C}$,
$\mathbb{P}(x_{t+T+1}\in A\ |\ x_{t}=x)$ $=\mathbb{P}(x_{T+1}\in A\ |\ x_{0}=x)$
$\geq\mathbb{P}(x_{T+1}\in A\ |\ x_{1}\in B,x_{0}=x)\cdot\mathbb{P}(x_{1}\in B\ |\ x_{0}=x)\geq\epsilon,$
@@
$\epsilon=\delta\cdot\inf_{y\in\mathcal{C}}\mathbb{P}(x_{1}\in B\ |\ x_{0}=y)>0,$
@@
This verifies that the chain $\{x_{t}\}$ is Harris recurrent (see Section 9 in [23] for definition). It can be further derived that for any $x\in\mathcal{C}$,
@@
The bound above is uniform for all $x\in\mathcal{C}$ and this implies that $\mathcal{C}$ is a regular set of $\{x_{t}\}$ (see Section 11 in [23] for definition). Finally, one can apply Theorem 13.0.1 in [23] to conclude that there exists a unique invariant probability measure $\nu_{\mathcal{D}}$ on $\mathcal{B}(\mathcal{C})$ and that the distribution of $x_{t}$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ conditioned on $x_{0}=x$ for any $x\in\mathcal{C}$. ∎
Appendix F Proof of Theorem 3.2
@@
Definition F.1 (Shifted Rényi divergence).
@@
Lemma F.2 (Data-processing inequality for Rényi divergence, Lemma 2.6 in [16]).
@@
$D_{\alpha}(h_{\#}\mu||h_{\#}\nu)\leq D_{\alpha}(\mu||\nu),$ (7)
where $h_{\#}\mu$ denotes the pushforward operation for a function $h$ and distribution $\mu$.
Proposition F.3 (Weak Triangle Inequality of Rényi divergence, Corollary 4 in [19]).
@@
Recall that from the sketch of proof of Theorem 3.2, we have defined the six distributions: $\nu_{\mathcal{D}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0},\rho_{K|\mathcal{B}}^{0}$ are the stationary distribution of the learning process, distribution at epoch $T$ of the learning process and distribution at epoch $K$ of the unlearning process. Note that we learn on dataset $\mathcal{D}$ and fine-tune on $\mathcal{D}^{\prime}$. On the other hand, the corresponding distributions of “adjacent” processes that learn on $\mathcal{D}^{\prime}$ and still unlearn on $\mathcal{D}^{\prime}$ are denoted as $\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime},\rho_{K|\mathcal{B}}^{0,\prime}$ similarly. Note that $\nu_{T|\mathcal{B}}^{0,\prime}$ is the distribution of retraining from scratch on $\mathcal{D}^{\prime}$, and we aim to bound $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})$ for all possible $\mathcal{D},\mathcal{D}^{\prime}$ pairs. By Proposition F.3, we know that for any $\alpha>1$, by choosing $p=q=2$, we have
@@
Recall the Definition 2.1 that $d_{\alpha}(P,Q)=\max(D_{\alpha}(P||Q),D_{\alpha}(Q||P))$ for distributions $P,Q$. The above inequality is correct since consider any distributions $P,Q,R$, by Proposition F.3 we have
@@
Finally, due to data-processing inequality for Rényi divergence (Lemma F.2), applying the final projection step does not increase the corresponding Rényi divergence. As a result, by Lemma 2.6 we have
@@
Appendix G Proof of Theorem 3.11
@@
The proof is a direct application of triangle inequality, Lemma 3.4 and 3.3. We will prove it by induction. For the base case $s=1$ it trivial, since $\bar{\nu}_{\mathcal{D}_{0}}=\nu_{\mathcal{D}_{0}}$ as we choose $y_{0}^{0,(0)}=x_{\infty}^{0}$. Thus by our definition that $Z_{\mathcal{B}}$ is the upper bound of $W_{\infty}(\nu_{\mathcal{D|\mathcal{B}}},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ for any adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$. Apparently, we also have
@@
Appendix H Proof of Corollary 3.7
@@
Appendix I Improved bound with randomized $\mathcal{B}$
Corollary.
where $Z_{\mathcal{B}}$ is the bound described in Lemma 3.3.
We first restate the lemma in [22], which is an application of the joint convexity of KL divergence.
Lemma I.1 (Lemma 4.1 in [22]).
@@ Appendix I Improved bound with randomized $\mathcal{B}$
The proof of Corollary 3.8 directly follows by repeating the proof of Lemma 3.7 but with the $\mathcal{B}$ dependent $Z_{\mathcal{B}}$ described in Lemma 3.3. Then leverage Lemma I.1 leads to the result.
Appendix J $W_{\infty}$ bound for batch unlearning
Corollary J.1 ($W_{\infty}$ bound for batch unlearning).
$W_{\infty}(\nu_{T|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})\leq\min\left(\frac{1-c^{T}}{1-c}\sum_{g=0}^{G-1}c^{n/b-j_{g}-1}\frac{2\eta MS_{j_{g}}}{b},2R\right).$
@@
The proof is a direct generalization of Lemma 3.3. Recall that for the per-iteration bound within an epoch, there are two possible cases: 1) we encounter the mini-batch $\mathcal{B}^{j}$ that contains indices of replaced points 2) or not. This is equivalently to 1) $j\in\{j_{g}\}_{g=0}^{G-1}$ or 2) $j\notin\{j_{g}\}_{g=0}^{G-1}$.
@@
$\|\psi_{j}(x_{t+1}^{0})-\psi_{j}^{\prime}(x_{t+1}^{0,\prime})\|\leq(1-\eta m)^{n/b}\|\psi_{j}(x_{t}^{0})-\psi_{j}^{\prime}(x_{t}^{0,\prime})\|+\sum_{g=0}^{G-1}(1-\eta m)^{n/b-j_{g}-1}\frac{2\eta MS_{j_{g}}}{b}.$ (43)
The rest analysis is the same as those in Lemma 3.3, where we iterate over $T$ epochs, infimum over all possible coupling, and then simplify the expression using geometric series. Thus we complete the proof.
Appendix K Proof of Proposition 3.9
Proposition.
Under the same setting as Corollary 3.7 with $b=n$, $\eta\leq\frac{m}{2L^{2}}$ and assume the minimizer of any $f_{\mathcal{D}}$ is in the relative interior of $\mathcal{C}_{R}\subseteq\mathbb{R}^{d}$, for any given adjacent dataset pair $\mathcal{D},\mathcal{D}^{\prime}$ the output of the $K^{th}$ unlearning iteration $y_{K}^{0}$ satisfies
$\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0})-\inf_{x\in\mathcal{C}_{R}}f_{\mathcal{D}^{\prime}}(x)]\leq Mc^{K}\min(\frac{1}{1-c}\frac{2\eta M}{n},2R)+\frac{2Ld\sigma^{2}}{m}.$ (44)
The proof is based on the following key lemma from [24].
Lemma K.1.
For $M$-Lipschitz, $m$-strongly convex and $L$-smooth loss function $f_{\mathcal{D}}(x)$ over the closed bounded convex set $\mathcal{C}_{R}\subseteq\mathbb{R}^{d}$ with radius $R$, step size $\eta\leq\frac{m}{2L^{2}}$ and initial parameter $x_{0}\sim\nu_{0}$, the excess empirical risk of the learning process of Algorithm 1 for $T$ epoch is bounded by
$\mathbb{E}[f_{\mathcal{D}}(x_{T}^{0})-f_{\mathcal{D}}(x^{\star})]\leq\frac{L}{2}\exp(-m\eta T)\mathbb{E}_{x_{0}\sim\nu_{0}}[\|x_{0}-x^{\star}\|^{2}]+\frac{2Ld\sigma^{2}}{m},$ (45)
where $x^{\star}$ is the minimizer of $f_{\mathcal{D}}(x)$ in the relative interior of $\mathcal{C}_{R}$ and $d$ is the dimension of parameter.
Clearly, we can see that under the convergent assumption of Corollary 3.7, the excess risk of $y_{0}^{0}=x_{\infty}$ is
$\mathbb{E}[f_{\mathcal{D}}(y_{0}^{0})-f_{\mathcal{D}}(x^{\star})]\leq\frac{2Ld\sigma^{2}}{m}.$ (46)
This is true since $\|x_{0}-x^{\star}\|\leq 2R$ is finite by the boundedness of $\mathcal{C}_{R}$.
Now we are ready to prove Proposition 3.9.
Proof.
Start by observing that
$\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0})-f_{\mathcal{D}^{\prime}}(y^{\star})]=\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0})-f_{\mathcal{D}^{\prime}}(y_{K}^{0,\prime})]+\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0,\prime})-f_{\mathcal{D}^{\prime}}(y^{\star})],$ (47)
where we recall that $y_{K}^{0,\prime}$ is the “adjacent process” of $y_{K}^{0}$ that running noisy gradient descent on $\mathcal{D}^{\prime}$ for both learning and unlearning processes. As a result, we have $y_{K}^{0,\prime}=x_{T+K}^{0,\prime}$ for $T$ learning epoch. By taking $T\rightarrow\infty$ under the convergent assumption and utilizing Lemma K.1, we have
$\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0,\prime})-f_{\mathcal{D}^{\prime}}(y^{\star})]\leq\frac{2Ld\sigma^{2}}{m}.$ (48)
Hence we are left with analyzing the first term. We will bound it in terms of $\|y_{K}^{0}-y_{K}^{0,\prime}\|$. Note that we choose a particular coupling between $y_{K}^{0},y_{K}^{0,\prime}$ as in Lemma 3.3, where the Gaussian noise in $y_{K}^{0},y_{K}^{0,\prime}$ are identical almost surely. Also for simplicity, we drop 00 in the superscript since we are analyzing the full batch case. By $M$-Lipschitzness of $f_{\mathcal{D}^{\prime}}$, we have
$\|f_{\mathcal{D}^{\prime}}(y_{K})-f_{\mathcal{D}^{\prime}}(y_{K}^{\prime})\|\leq M\|y_{K}^{0}-y_{K}^{0,\prime}\|.$ (49)
For $\|y_{K}^{0}-y_{K}^{0,\prime}\|$, we have
$\|y_{K}-y_{K}^{\prime}\|$ (50)
$=\|\Pi_{\mathcal{C}_{R}}[y_{K-1}-\eta\nabla f_{\mathcal{D}^{\prime}}(y_{K-1})+\sqrt{2\eta\sigma^{2}}W_{K-1}]-\Pi_{\mathcal{C}_{R}}[y_{K-1}^{\prime}-\eta\nabla f_{\mathcal{D}^{\prime}}(y_{K-1}^{\prime})+\sqrt{2\eta\sigma^{2}}W_{K-1}]\|$ (51)
$\leq\|y_{K-1}-\eta\nabla f_{\mathcal{D}^{\prime}}(y_{K-1})-(y_{K-1}^{\prime}-\eta\nabla f_{\mathcal{D}^{\prime}}(y_{K-1}^{\prime}))\|$ (52)
$\leq(1-\eta m)\|y_{K-1}-y_{K-1}^{\prime}\|,$ (53)
where the first inequality is due to the fact that $\Pi_{\mathcal{C}_{R}}$ is a contraction map and noise cancels out due to our coupling of $y_{K}^{0},y_{K}^{0,\prime}$. The second inequality is due to the fact that gradient update is $(1-m\eta)$-contraction map when $f_{\mathcal{D}^{\prime}}$ is $L$-smooth, $m$-strongly convex and $\eta\leq\frac{1}{L}$. By iterating this bound, we have
$\|f_{\mathcal{D}^{\prime}}(y_{K})-f_{\mathcal{D}^{\prime}}(y_{K}^{\prime})\|\leq M(1-m\eta)^{K}\|y_{0}-y_{0}^{\prime}\|$ (54)
$\leq Mc^{K}\min(\frac{1}{1-c}\frac{2\eta M}{n},2R),$ (55)
where $c=1-m\eta$ and the last inequality is due to Lemma 3.3 for the case $b=n$ and taking the limit $T\rightarrow\infty$. Combining all results together, we have
$\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0})-f_{\mathcal{D}^{\prime}}(y^{\star})]$ (56)
$\leq\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0})-f_{\mathcal{D}^{\prime}}(y_{K}^{0,\prime})]+\frac{2Ld\sigma^{2}}{m}$ (57)
$\leq\mathbb{E}[\|f_{\mathcal{D}^{\prime}}(y_{K}^{0})-f_{\mathcal{D}^{\prime}}(y_{K}^{0,\prime})\|]+\frac{2Ld\sigma^{2}}{m}$ (58)
$\leq Mc^{K}\min(\frac{1}{1-c}\frac{2\eta M}{n},2R)+\frac{2Ld\sigma^{2}}{m}.$ (59)
Together we complete the proof. ∎
Appendix L Proof of Lemma 3.3
@@
We prove the bound of one iteration and iterate over that result under a specific coupling of the adjacent PNSGD learning processes, which is by definition an upper bound for taking infimum over all possible coupling. Given an mini-batch sequence $\mathcal{B}$, assume the adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$ differ at index $i_{0}\in[n]$ and $i_{0}$ belongs to $j_{0}^{th}$ mini-batch (i.e., $i_{0}\in\mathcal{B}^{j_{0}}$. For simplicity, we drop the condition on $\mathcal{B}$ for all quantities in the proof as long as it is clear that we always condition on a fixed $\mathcal{B}$. Let us denote $\psi_{j}(x)=x-\frac{\eta}{b}\sum_{i\in\mathbf{B}^{j}}\nabla f(x;\mathbf{d}_{i}^{\prime})$ and similar for $\psi_{j}^{\prime}$ on $\mathcal{D}^{\prime}$. Note that for some coupling of $\nu_{t}^{j},\nu_{t}^{j,\prime}$ denoted as $\gamma_{t}^{j}$,
@@
Now, note that by definition the noise $W_{t}^{j}$ is independent of $x_{t}^{j}$, we can simply choose the coupling $\gamma$ so that $W_{t}^{j}=W_{t}^{j,\prime}$ for all $t,j$ and $x_{0}^{0}=x_{0}^{0,\prime}$ due to the same initialization distribution. So the last term is 00. For the first term, there are two possible cases: 1) $j=j_{0}$ and 2) $j\neq j_{0}$. For case 1), we have
@@
Combining these two cases, we have
$\text{esssup}_{\gamma}\|x_{t}^{j+1}-x_{t}^{j+1,\prime}\|\leq\text{esssup}_{\gamma}(1-\eta m)\|x_{t}^{j}-x_{t}^{j,\prime}\|+\frac{2\eta M}{b}\mathbf{1}[j=j_{0}]$ (73)
@@
$\text{esssup}_{\gamma}\|x_{t+1}^{0}-x_{t+1}^{0,\prime}\|\leq\text{esssup}_{\gamma}(1-\eta m)^{n/b}\|x_{t}^{0}-x_{t}^{0,\prime}\|+(1-\eta m)^{n/b-j_{0}-1}\frac{2\eta M}{b}$ (75)
@@
$\text{esssup}_{\gamma}\|x_{T}^{0}-x_{T}^{0,\prime}\|\leq\text{esssup}_{\gamma}(1-\eta m)^{Tn/b}\|x_{0}^{0}-x_{0}^{0,\prime}\|+\sum_{t=0}^{T-1}(1-\eta m)^{(t+1)n/b-j_{0}-1}\frac{2\eta M}{b}$ (77)
@@
where the last equality follows by our choice of coupling $\gamma$ such that $x_{0}^{0}=x_{0}^{0,\prime}$ due to the same initialization distribution. Now by taking the infimum overall possible coupling $\gamma$ and simply the expression, we have
@@
Appendix M Proof of Lemma 3.4